If the man’s age is a 3-digit number, his age is equal to product of two numbers. One of the numbers is the sum of the digits and the other number is product of the digits his age is?
Answer by ikleyn(52788)

.


One possible solution is the three-digits number  144.

The sum of its digits is   1 + 4 + 4 = 9;
the product of its digits is   1*4*4 = 16.

And the product of the sum of the digits by the product of the digits is   9*16 = 144.


E n j o y (!)